OpenStudy (anonymous):

ok so You roll one number cube twice. Find each probability. 1: P(1 and odd)

OpenStudy (anonymous):

How do i do this?

OpenStudy (anonymous):

First find the probability of P(1), which is 1/6 Then find P(odd), which is 3/6 = 1/2

OpenStudy (anonymous):

But how did u get 3/6?

OpenStudy (anonymous):

Multiply the two probabilities together to find the probability they will happen in sequence. 1/6 * 1/2 = 1/12

OpenStudy (anonymous):

Wait im Confused what am i Suppose to Rigth?

OpenStudy (anonymous):

I found 3/6 because there are 3 odd numbers out of 6 total

OpenStudy (anonymous):

and 1/6 because there is one 1 and 6 total numbers on a die

OpenStudy (anonymous):

the final answer is 1/12

OpenStudy (anonymous):

so i write to answer it P(1/6x1/2=1/12?

OpenStudy (anonymous):

P(1) = 1/6 P(odd) = 1/2 P(1)*P(odd) = 1/12

OpenStudy (anonymous):

Can uAlso Help me with this P(3 and even is the even 2?